# Let’s Influence Algorithms Together: How Millions of Fans Build Collective Understanding of Algorithms and Organize Coordinated Algorithmic Actions

### v1

@@ 3.3.1. Online and on-site observation
We joined core fan groups on various social media platforms, such as WeChat, Weibo, and WhatsApp, where algorithms for organizing and coordinating actions are discussed, to understand how core fans perceive algorithms and how they extend this understanding to collective actions. Additionally, the two lead researchers participated in 22 offline gatherings of core fans to better grasp their perspectives and actions. We also joined the social media groups of general fans, focusing on discussions surrounding folk understanding of algorithms and the coordination of collective algorithmic actions. In these instances, the researchers were not passive participants; rather, they engaged in participatory observation as part of the fan community, fully immersing themselves in these algorithm-related fan activities (Zheng and Xiao, 2023). This participatory observation approach follows an insider’s perspective, allowing the two lead authors to document and discuss the activities of core fans and the virtual or physical spaces they visited, thereby enabling interactive verification of data collection with other interviewees (Zheng and Xiao, 2023). These observations spanned from 2021 to 2023.
@@ 3.3.2. Interviews with core fans and general fans
Semi-structured interviews often effectively complement ethnographic observation data (Thach et al., 2023). Therefore, we conducted semi-structured interviews with the 43 core fans we observed and an additional 15 general fans from their fan clubs. These interviews were conducted online following the ethnographic observations to gather their views on collective algorithmic actions. All interviews were recorded and transcribed by researchers. The semi-structured interviews with core fans focused on their motivations for organizing collective algorithmic actions, challenges, the distinctions between different algorithmic actions, the iteration processes, and the reasoning behind these iterations. These interviews for core fans ranged from 1 to 3 hours, depending on the core fans’ availability. For the general fans, the semi-structured interviews explored their motivations for participating in collective algorithmic actions, their perceptions of the effectiveness of these actions, their views on the persuasive rhetoric and algorithmic action tutorials organized by core fans, and their assessment of the overall performance of the fan communities and fan clubs. These interviews for general fans ranged from 1 to 1.5 hours.
3.3.3. Focus group with core fans
After conducting observations and interviews, we held three focus group sessions, each with an average of five participants lasting approximately two and a half hours, to gather a collective perspective on algorithmic understanding and actions among core fans. In HCI research, focus groups are a common complement to ethnography, enabling the exploration of group dynamics, collective reasoning, and the communal construction of knowledge. Unlike individual interviews, focus groups facilitate interactions among participants, bringing to light shared concerns, common strategies, and differing viewpoints—all crucial for understanding the collective behaviors that drive fan-led algorithmic practices.
@@ 3.4. Data Analysis
Through a thematic analysis approach (Braun and Clarke, 2012, 2019), two researchers conducted a comprehensive qualitative analysis of the field notes, interview transcripts, and focus group discussions. By applying open coding(Braun and Clarke, 2012, 2019), the researchers could break down the data into discrete elements, making it possible to identify themes and relationships that might not have been immediately apparent. This collaborative process of discussion, reflection, and refinement also helped to enhance the reliability of the analysis. We did not calculate inter-coder reliability between the two researchers. Instead, we adhered to the principles of reflexive thematic analysis(Braun and Clarke, 2012, 2019), engaging in ongoing critical reflection and open discussions to address any consensus or discrepancies that arose during the coding process. After completing the coding, the research team conceptualizes higher-level themes from these codes, employing the affinity diagram method to organize the emerging themes. This process ultimately resulted in the identification of first-level themes, second-level themes, and third-level themes. Our findings are detailed in Section 4, with the section headings aligning closely with the second and third-level themes. A comprehensive list of all second and third-level themes can be found in the appendix ( Table 1).
@@ 3.5. Ethics
4. Results
Algorithmic labor is a prevalent phenomenon among fan communities, primarily involving fans engaging in intentional actions through platform algorithms to enhance the visibility of their idols. This includes increasing repost numbers on social media platforms, video platform views, and music platform rankings. These actions are often repetitive, named “algorithmic labor” in fandom (Zheng and Xiao, 2023; Zhang et al., 2023). In Chinese fandom, fans sometimes refer to this behavior as ”data labor” because they are directly engaged with various forms of data, such as shares and views, while the algorithm remains invisible.
4.1. “All for Our Idol”: Recognizing the Necessity of Collective Algorithmic Action (RQ1)
@@ 4. Results
The first step in fan-driven collective algorithmic action is recognizing the necessity of such action. Core fans and general fans, when asked why fans engage in algorithmic action, unanimously see it as a key way to support their idols. They clearly understand that the behavior of fans on social media has a significant impact on the visibility of their idols, especially when influenced by recommendation algorithms. They have made a conscious connection between how fans engage with social media and the level of exposure their idols receive. By closely examining the various types of data generated by fan activities, they have been able to distinguish how different forms of social media engagement—such as likes, shares, comments, and other interactions—affect the visibility and prominence of their idols within the algorithmically driven visibility of social media platforms.
@@
In the view of the core fans we interviewed, recognizing the impact of fan-platform interactions on their idols, as well as understanding the specific roles of different forms of interaction, is the first step in engaging in algorithmic action. This awareness makes them realize that taking algorithmic action is necessary and feasible. CF43 explains to us: ”First, we need to understand that developing algorithms is beneficial for idols, right? And this benefit is quite significant; otherwise, we wouldn’t be working so hard on this. Second, we must be clear about the practical impact of our approach; we can’t just run around like headless chickens.” Next, we will explore how these fans perceive and interpret various fan behaviors on social media.
4.1.1. The Importance of Data Performance in Idols’ Commercial Collaborations
@@
Before organizing collective algorithmic actions, people must first recognize that such actions are essential and necessary. Fans believe that an idol’s data performance on social media significantly influences their commercial opportunities. CF38, a fan club leader, explains that brands don’t just subjectively choose celebrities as ambassadors; objective data on social media is crucial. Brands work with data analysis firms and social media or e-commerce platforms for relevant data. This data significantly influences their choice of an idol.
@@
CF14, a fan club administrator, said that an idol partnering with a top brand gains not only significant monetary benefits but also promotional support from the brand. This includes invitations to events and shows sponsored by the brand.
@@
Certainly, impressive ”data performance” is not limited to securing brand endorsements. CF42, a devoted fan with 15 years of fan experience, illustrated this point with one of her 2021 posts on Weibo, which stated that not only commercial brands but also reality shows and galas consider ”data performance” when selecting their guests. While these platforms might not always seek assistance from professional data companies, given their less commercial nature, they still rely on intuitive and direct data metrics as significant factors in choosing celebrities as guests. Finally, this post garnered over 5,000 likes, 600 reposts, and 500 comments on Weibo, earning widespread recognition from the fan community.
@@
4.1.2. Perceived Relationship Between Fan-Driven Collective Algorithmic Action and Social Media Algorithms
@@
After recognizing that such action is necessary, people must also understand that their efforts, especially large-scale collective actions, can indeed influence algorithms in their perception. In the views of fans, their actions on social media are likely to influence, or even manipulate, the algorithms of these platforms.
@@
Fans first distinguish between data that can be achieved without manipulating the algorithm and data performance that requires algorithmic intervention, thereby establishing the necessity for fans to take action against the algorithm. CF29, a core fan responsible for organizing fan-driven data performances, outlined the three primary dimensions of core data: (1) monetary-related data, (2) data pertaining to votes and rankings, and (3) social media engagement data.
The monetary data involves sales volumes linked to the idol, encompassing album sales, magazine sales, and endorsed product sales. This data typically reflects the commercial influence of the idol’s core fanbase, indicating their purchasing power and the potential for a brand to recoup costs from idol endorsements. The data pertaining to votes and rankings gauge the willingness of the idol’s core fanbase to devote time and effort, potentially acting as a ’water army’—a term for those hired to manipulate online content—for the brand without compensation.
CF29 noted, ”When our idol is chosen as a brand ambassador, our data performance efforts indirectly benefit the brand, as we engage in activities like reposting, liking, and commenting on specific posts or videos.” She added, ”An idol’s social media engagement data is closely tied to their public visibility and reputation. Typically, better engagement data equates to stronger publicity.” CF33, an administrator of a data team in fandom, further explained that in the realm of reality shows, an idol’s public exposure is often seen as indicative of potential program ratings, making publicity data a critical factor in the invitation process for reality shows.
@@
The first two types of data, concerning monetary contributions and votes or rankings, gauge the engagement of an idol’s core fans. Fans can elevate their idol’s data through straightforward methods like repeated voting and financial expenditure. However, the third type of data, social media engagement data, cannot be manipulated directly. This is because social media platforms employ complex algorithms to filter out artificial publicity data and focus on acquiring more genuine engagements. Therefore, to enhance this type of data performance, fans must carefully devise strategies to influence the algorithm.
CF29 provided an example regarding the order arrangement within comments on Weibo. According to CF29, Weibo utilizes a metric termed ”heat” for its specific algorithm to rank comments. While ”heat” is not directly visible to users, it plays a crucial role in how the platform orders comments. The exact components of ”heat” are not publicly known.
Based on fans’ observations, they believe it encompasses the number of likes, the volume and likes of nested comments, the commenter’s account quality, and user interaction. Fans liken the dense layers of comments, which can number in the hundreds or thousands, to a ’building,’ with nested comments (comments on comments) being a ’building within a building.’ However, Weibo does not disclose the precise criteria for calculating ”heat.”
This lack of transparency arises because fans often strategically like or comment on posts favorable to their idols to boost those comments’ ”heat,” thereby increasing their visibility. This tactic, known as ”controlling comments under posts,” which means conscious control of the comments section by fans, is seen by the public as significantly impacting their online experience, leading to calls for Weibo to mitigate the influence of fan-driven ”controlling comments.” Consequently, Weibo has revised its rules for ”heat” multiple times. Despite these changes, fans continually adapt, tweaking their strategies to align with the new algorithmic logic. This dynamic results in a delicate balance between the platform’s efforts to maintain authentic engagement and the fans’ endeavors to promote their idols.
Fans recognize that their various actions across multiple platforms can significantly impact their idol’s digital visibility, prompting them to continuously expand the scope of their actions. In addition to ”controlling comments” beneath Weibo posts, several other digital arenas have emerged as critical battlegrounds in the interaction between fans and algorithms. These include trending topics on social media, popularity rankings on video platforms, view counts, and algorithmic recommendations. Each area represents a significant opportunity for fans to influence the online visibility and perceived popularity of their favorite celebrities. By engaging with these platforms, fans attempt to navigate and sometimes manipulate algorithmic systems to boost their idol’s digital presence and success.
@@
4.2. Fan-Driven Folk Theories: The Foundation for Guiding Collective Algorithmic Actions (RQ2)

### v2

@@ 3.4.1. Online and on-site observation
We joined core fan groups on various social media platforms, such as WeChat, Weibo, and WhatsApp, where algorithms for organizing and coordinating actions are discussed, to understand how core fans perceive algorithms and how they extend this understanding to collective actions. Additionally, the two lead researchers participated in 22 offline gatherings of core fans to better grasp their perspectives and actions. We also joined the social media groups of general fans, focusing on discussions surrounding folk understanding of algorithms and the coordination of collective algorithmic actions. In these instances, the researchers were not passive participants; rather, they engaged in participatory observation as part of the fan community, fully immersing themselves in these algorithm-related fan activities (Zheng and Xiao, 2023).
In most cases, we observed the internal discussions among core fans and the interactions between general fans on a daily basis across the 12 fan clubs. These discussions typically took place within different dedicated chat groups, reflecting the segmented nature of communication within fan communities. In accordance with the participants’ requests, we did not directly collect chat records using methods such as screenshots, but instead used secondary field notes.
This participatory observation approach follows an insider’s perspective, allowing the two lead authors to document and discuss the activities of core fans and the virtual or physical spaces they visited, thereby enabling interactive verification of data collection with other interviewees (Zheng and Xiao, 2023). Data collection during online observations included taking detailed field notes of the types of content shared, the language used in organizing algorithmic actions, and the dynamics of group interactions. During on-site observations at offline gatherings, the researchers recorded their observations in field journals immediately after the events to ensure accuracy and minimize recall bias. In both settings, the researchers adhered to ethical principles of informed consent and ensured that their presence did not disrupt the natural flow of fan activities. These observations spanned from 2021 to 2023.
Throughout these two years, as the part of ethnography, we also engaged with Chinese core fans biweekly to gather their reflections on their daily collective algorithmic actions. These informal reflections were not structured as formal semi-structured interviews but were instead conducted in a conversational manner, resembling casual discussions. This provided a relaxed environment where participants felt comfortable sharing insights about both the internal strategies within the core fan groups and their influence on general fan groups. This approach aligns with ethnographic principles that emphasize immersion in the everyday lives of participants and the natural flow of information exchange. Regular interactions over an extended period enabled us to build trust with participants, providing a deeper understanding of the shifting dynamics within fan communities and the evolving nature of algorithmic practices. It also allowed us to document not just isolated actions but the iterative processes and ongoing adaptations that underpin successful collective actions.
@@ 3.4.2. Interviews with core fans and general fans
Semi-structured interviews often effectively complement ethnographic observation data (Thach et al., 2023; Schensul et al., 1999). Therefore, we conducted semi-structured interviews with the 43 core fans we observed and an additional 15 general fans from their fan clubs. These interviews were conducted online following the ethnographic observations to gather their views on collective algorithmic actions. All interviews were recorded and transcribed by the researchers. The semi-structured interviews with core fans included open-ended questions such as: ”What motivates you to organize algorithmic actions?” ”How do you perceive the effectiveness of these actions?” and ”Can you describe the process of designing persuasive messages for general fans?” For general fans, questions included: ”What motivates you to participate in algorithmic actions?” ”How do you perceive the instructions provided by core fans?” and ”What challenges, if any, have you faced in engaging with these actions?” These interview questions were piloted with a small group of fans to ensure clarity and relevance.
The semi-structured interviews with core fans focused on their motivations for organizing collective algorithmic actions, challenges, the distinctions between different algorithmic actions, the iteration processes, and the reasoning behind these iterations. These interviews for core fans ranged from 1 to 3 hours, depending on the core fans’ availability. For general fans, the semi-structured interviews explored their motivations for participating in collective algorithmic actions, their perceptions of the effectiveness of these actions, their views on the persuasive rhetoric and algorithmic action tutorials organized by core fans, and their assessment of the overall performance of the fan communities and fan clubs. These interviews for general fans ranged from 1 to 1.5 hours.
3.4.3. Focus groups with core fans
After conducting observations and interviews, we held three focus group sessions with core fans, each with an average of five participants lasting approximately two and a half hours, to gather a collective perspective on algorithmic understanding and actions among core fans. Each focus group session includes five different participants. The focus group sessions were guided by a researcher moderator, who posed open-ended prompts such as: ”What are the key challenges in organizing collective algorithmic actions?” and ”How do you as a group decide on strategies for engaging general fans?” The moderator ensured that all participants had the opportunity to contribute while maintaining the focus of the discussion. These sessions were recorded with participants’ consent, and transcripts were analyzed to identify themes related to group decision-making, shared challenges, and communal learning processes.
In HCI research, focus groups are a common complement to ethnography, enabling the exploration of group dynamics, collective reasoning, and the communal construction of knowledge. Unlike individual interviews, focus groups facilitate interactions among participants, bringing to light shared concerns, common strategies, and differing viewpoints, all crucial for understanding the collective behaviors that drive fan-led algorithmic practices.
@@ 3.5. Data Analysis
In this study, we employed a mixed-methods ethnographic approach, integrating field notes, interview transcripts, and focus group discussions to capture the multifaceted dynamics of our research context. Following established practices in ethnographic research, such as those outlined by Braun and Clarke in their thematic analysis framework (Braun and Clarke, 2012, 2019), we analyzed all data collectively without segregating it by source or method. This reflects the ethnographic tradition of treating diverse data as complementary pieces of a holistic puzzle, where multiple methods serve to triangulate findings and deepen understanding rather than isolating specific insights to a single methodological tool.
Our data analysis followed a reflexive thematic analysis approach (Braun and Clarke, 2012, 2019), in which two researchers collaboratively conducted qualitative coding. The coding process began with both researchers independently coding a subset of data to familiarize themselves with the content and develop initial impressions. After this initial phase, the researchers convened to discuss their findings, consolidate preliminary codes, and create a unified codebook. This iterative process ensured alignment and consistency in coding while maintaining space for diverse interpretations. Separate codebooks were not created for the different data sources (observational data, interviews, focus groups), as our holistic approach treated these as interconnected data streams that contributed to a comprehensive understanding of the phenomenon. Instead, all data were coded using the same unified codebook, which evolved dynamically as additional insights emerged.
Through open coding, the data was systematically broken down into discrete elements, enabling the identification of underlying patterns, themes, and relationships. This step ensured that no potentially significant details were overlooked, as the researchers continuously revisited the data to refine their interpretations. Reflexive thematic analysis diverges from rigid, quantifiable measures of inter-coder reliability; instead, it emphasizes ongoing critical reflection and collaborative dialogue to ensure that diverse perspectives are considered throughout the analysis process.
To synthesize and organize our insights, we used the affinity diagram method (Harboe and Huang, 2015), a widely recognized tool for clustering qualitative data into hierarchical themes. The coding process resulted in the identification of a series of first-level, second-level, and third-level themes, each representing a progressively deeper abstraction of the data. This hierarchical organization not only helped to structure our findings but also allowed us to trace connections between granular observations and overarching theoretical concepts. The complete list of second- and third-level themes is provided in the Table 2 in appendix.
By analyzing data holistically and integrating diverse methodological tools, we adhered to the ethnographic principle of contextual immersion. This approach emphasizes understanding the research phenomenon as a dynamic interplay of participant actions, cultural norms, and systemic influences. It also acknowledges the value of participant narratives, enabling us to co-construct meanings with our subjects rather than imposing rigid analytical frameworks (Braun and Clarke, 2012, 2019). This integrated and reflexive approach allowed us to derive rich, contextually grounded insights that form the foundation of our subsequent analysis. We also present our research questions, corresponding findings, and their respective sections in Table 1 to enhance the clarity.
3.6. Positionality Statement
We recognize that our personal experiences and backgrounds have shaped our research approach. Our team members bring diverse expertise in HCI, Communication and Sociology, with some research experience in folk theories of algorithms and fandom studies. Two authors have been actively involved in fan-driven algorithmic practices for over five years. One author, in particular, is an active core fan within her fan club, managing her own account in fandom that boasts nearly twenty thousand followers on Weibo and an average monthly readership of thirty million. She has extensive experience organizing fan-driven algorithmic actions across multiple countries, including China, South Korea, Malaysia, and Indonesia. Two authors have extensive experience in fan studies, while the other focus on algorithm research. In our research, the team members who are part of the fan community continuously emphasized the importance of adhering to community norms and ethical considerations. Meanwhile, the other researchers, serving as outsiders, provided a critical perspective. We regularly organized internal discussions among the researchers to ensure ethical considerations were upheld.
@@ 3.7. Ethics
| Research Question (RQ) | Key Findings and Explanations |
| RQ1: What motivates fans to mobilize collective algorithmic actions? | section 4: Fans recognize the necessity of collective algorithmic action, believing their idol’s visibility and success depend on collective efforts, often summarized as ”all for our idol.” |
| RQ2: How do fans build a collective folk understanding of algorithms? | section 5: Fans develop a collective folk understanding of algorithms, providing informal insights into how algorithms work across platforms and countries, forming the foundation for effective actions. |
| RQ3: How do fans organize and execute collective algorithmic actions? | section 6 and section 7: Rational persuasion alone is insufficient. Effective actions rely on shared emotionality in fan communities, encapsulated in sentiments like ”our idol only has us.” Core fans must motivate general fans and use their folk understanding of algorithms to drive participation. |
4. Findings: “All for Our Idol”: Recognizing the Necessity of Collective Algorithmic Action (RQ1)
@@
The first step in fan-driven collective algorithmic action is recognizing the necessity of such action. Core fans and general fans, when asked why fans engage in algorithmic action, unanimously see it as a key way to support their idols. They clearly understand that the behavior of fans on social media has a critical impact on the visibility of their idols, especially when influenced by recommendation algorithms. They have made a conscious connection between how fans engage with social media and the level of exposure their idols receive. By closely examining the various types of data generated by fan activities, they have been able to distinguish how different forms of social media engagement, such as likes, shares, comments, and other interactions affect the visibility and prominence of their idols within the algorithmically driven visibility of social media platforms.
@@
In the view of the core fans we interviewed, recognizing the impact of fan-platform interactions on their idols, as well as understanding the specific roles of different forms of interaction, is the first step in engaging in algorithmic action. This awareness makes them realize that taking algorithmic action is necessary and feasible. CF43 explains to us: ”First, we need to understand that developing algorithms is beneficial for idols, right? And this benefit is quite useful; otherwise, we wouldn’t be working so hard on this. Second, we must be clear about the practical impact of our approach; we can’t just run around like headless chickens.” Next, we will explore how these fans perceive and interpret various fan behaviors on social media.
4.1. The Importance of Data Performance in Idols’ Commercial Collaborations
@@
Before organizing collective algorithmic actions, fans must first recognize that such actions are essential and necessary. Fans believe that an idol’s data performance on social media crucially influences their commercial opportunities. CF38, a fan club leader, explains that brands don’t just subjectively choose celebrities as ambassadors; objective data on social media is crucial. Brands work with data analysis firms and social media or e-commerce platforms for relevant data. This data meaningfully influences their choice of an idol.
@@
CF14, a fan club administrator, said that an idol partnering with a top brand gains not only notable monetary benefits but also promotional support from the brand. This includes invitations to events and shows sponsored by the brand.
@@
Certainly, impressive ”data performance” is not limited to securing brand endorsements. CF42, a devoted fan with 15 years of fan experience, illustrated this point with one of her 2021 posts on Weibo, which stated that not only commercial brands but also reality shows and galas consider ”data performance” when selecting their guests. While these platforms might not always seek assistance from professional data companies, given their less commercial nature, they still rely on intuitive and direct data metrics as iimportant factors in choosing celebrities as guests. Finally, this post garnered over 5,000 likes, 600 reposts, and 500 comments on Weibo, earning widespread recognition from the fan community.
@@
4.2. Perceived Relationship Between Fan-Driven Collective Algorithmic Action and Social Media Algorithms
@@
After recognizing the necessity of such collective actions, fans must also understand that their efforts, particularly large-scale collective actions, can indeed influence how algorithms perceive them. From the fans’ perspective, their activities on social media have the potential to affect and even manipulate these platforms’ algorithms.
@@
Fans begin by distinguishing between data that does not require algorithmic intervention and data that does. This step establishes the necessity of taking action against the algorithms. CF29, a core fan responsible for organizing fan-driven data performance, outlined three main types of critical data: (1) monetary data, (2) ranking-related data, and (3) social media engagement data.
Monetary data includes sales figures related to idols, such as album sales, magazine sales, and endorsed product sales. This data often reflects the commercial influence of an idol’s core fan base, showcasing their purchasing power and the potential for brands to recoup costs through idol endorsements.
Ranking-related data measures the time and effort fans are willing to invest in voting and boosting their idol’s rank. Fans may even act as “water armies”, a term for individuals hired (but in this case, unpaid) to manipulate online content to support brands. CF29 noted, “When our idol is chosen as a brand ambassador, our data performance indirectly benefits the brand, as we engage in activities like sharing, liking, and commenting on specific posts or videos.” She added, “The interaction data on social media directly correlates with the idol’s public exposure and reputation. Usually, better interaction data means more impactful promotion.” CF33, an administrator of a fan data team, explained that in the realm of reality TV, an idol’s public visibility often serves as a predictor of a show’s potential ratings, making engagement data crucial in securing show invitations.
The first two types of data, monetary contributions and ranking-related data, reflect the involvement level of an idol’s core fan base. Fans can easily improve these data metrics through repetitive actions like voting and spending money.
@@
However, the third type, social media engagement data, is more complex to manipulate because platforms use advanced algorithms to filter out artificial activity and focus on authentic engagement. To enhance performance in this area, fans must carefully devise strategies to influence the algorithms.
CF29 provided an example involving the ranking of comments on Weibo. According to CF29, Weibo uses a metric called “heat” to rank comments. Although the exact components of “heat” are not disclosed, it plays a critical role in determining comment rankings. Fans have deduced through observation that “heat” likely includes factors such as the number of likes, nested comments, account quality, and user interactions. They describe the dense layers of nested comments as akin to “buildings within buildings,” with hundreds or thousands of replies stacking up. Despite Weibo’s lack of transparency regarding the calculation of “heat,” fans strategically like and reply to favorable comments to increase their “heat” and boost their visibility. This practice, known as “controlling the comment section,” has drawn public criticism for affecting user experience, prompting Weibo to revise its “heat” rules multiple times. Nevertheless, fans continuously adapt their strategies to align with new algorithmic logic, creating a dynamic balance between the platform’s effort to ensure genuine engagement and fans’ efforts to promote their idols.
In summary, the three main types of data, monetary data, ranking data, and social media engagement data, represent distinct dimensions of fan-driven data performance, reflecting varying levels of effort, strategy, and influence. Monetary Data demonstrates the purchasing power of an idol’s core fan base, as seen in sales of albums, magazines, and endorsed products, showcasing their direct commercial value to brands. Ranking Data highlights fans’ dedication through repetitive tasks like voting, improving idols’ positions in rankings to showcase their popularity to the public and investors. These rankings often lead to rewards like increased exposure or credibility from prestigious awards. Social Media Engagement Data focuses on influencing recommendation algorithms. Unlike ranking data, “heat” is a more abstract concept without a transparent calculation method, leaving fans to rely on trial and error to understand it. For example, fans may strategically like and comment on TikTok videos to increase their visibility, a tactic also commonly used by international K-pop fans. Algorithms on platforms like X trends, Billboard rankings, and iTunes rankings are similarly influenced by overseas fans’ practices. In addition to platform-driven recommendations, social media engagement metrics, such as likes, comments, and shares, have become essential indicators of an idol’s popularity and interaction levels. These metrics are often benchmarks for comparison among fan communities, further solidifying their importance in fan strategies and idols’ success metrics.
@@
5. Findings: Fan-Driven Folk Theories: The Foundation for Guiding Collective Algorithmic Actions (RQ2)